OpenStudy (anonymous):

ILL BECOME A FAN IF YOU HELP ME !! Which of the following would best represent a sine function with an amplitude of 7, a period of pi over 2, and a midline at y = 3?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: y = A*sin(Bx - C) + D is the general sine equation |A| is the amplitude T = 2pi/B is the period y = D is the equation of the midline

OpenStudy (anonymous):

so 7 sin 8x + 3

jimthompson5910 (jim_thompson5910):

what is the value of B in this case?

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

so the period is T = 2pi/B T = 2pi/8 T = pi/4 but we want the period to be pi/2 instead

jimthompson5910 (jim_thompson5910):

So the equation can't be y = 7*sin(8x) + 3

OpenStudy (anonymous):

7sin (pu/4x)+3.

OpenStudy (anonymous):

pi*

jimthompson5910 (jim_thompson5910):

what's B this time?

OpenStudy (anonymous):

pi over 4

jimthompson5910 (jim_thompson5910):

T = 2pi/B T = 2pi/(pi/4) T = 2pi*(4/pi) T = 8pi/pi T = 8 Again, we don't get pi/2 like we want to, so it can't be that equation either

OpenStudy (anonymous):

would you habe pi or no

jimthompson5910 (jim_thompson5910):

the given period we want is T = pi/2 T = 2pi/B pi/2 = 2pi/B Solve for B to get B = ???

OpenStudy (anonymous):

4

OpenStudy (anonymous):

it would be